# Bidirectional Autoregressive Diffusion Model for Dance Generation

### v1

@@ Abstract
Dance serves as a powerful medium for expressing human emotions, but the lifelike generation of dance is still a considerable challenge. Recently, diffusion models have showcased remarkable generative abilities across various domains. They hold promise for human motion generation due to their adaptable many-to-many nature. Nonetheless, current diffusion-based motion generation models often create entire motion sequences directly and unidirectionally, lacking focus on the motion with local and bidirectional enhancement. When choreographing high-quality dance movements, people need to take into account not only the musical context but also the nearby music-aligned dance motions. To authentically capture human behavior, we propose a Bidirectional Autoregressive Diffusion Model (BADM) for music-to-dance generation, where a bidirectional encoder is built to enforce that the generated dance is harmonious in both the forward and backward directions. To make the generated dance motion smoother, a local information decoder is built for local motion enhancement. The proposed framework is able to generate new motions based on the input conditions and nearby motions, which foresees individual motion slices iteratively and consolidates all predictions. To further refine the synchronicity between the generated dance and the beat, the beat information is incorporated as an input to generate better music-aligned dance movements. Experimental results demonstrate that the proposed model achieves state-of-the-art performance compared to existing unidirectional approaches on the prominent benchmark for music-to-dance generation.
@@ 1 Introduction
In response to the limitations inherent in current methodologies, we present the innovative bidirectional Autoregressive diffusion model (BADM). Our model introduces a novel strategy by breaking down the entire noise sequence into smaller, manageable slices. During the generation of each slice, our model considers the preceding dance sequences and forwarding noise distributions using a cross-attention layer in a bidirectional way. The motivation is that the dances generated in both the forward and backward directions should be as similar as possible. Then refined noise slices are sent into the decoder along with conditions iteratively. The output dance slices are then concatenated and refined by a local information decoder to ensure a cohesive and harmonious overall sequence from a local perspective.
@@
Figure 2: Our bidirectional Autoregressive diffusion model (BADM) employs a denoising mechanism to enhance dance sequences from time $t=T$ to $t=0$. BADM begins with a noisy sequence $z_{T}$ at time $T$, and proceeds to generate an estimated dance sequence $\hat{x}$. The denoising procedure is iteratively applied until $t=0$. Our autoregressive (AR) model treats the whole noise sequence as $K$ slices. On the left, we show our model processing the $k$-th slice at the diffusion timestep $t$.
@@ 2.2 Music-to-dance Generation
The challenge of generating dances that synchronize with the input music has fueled the development of a wide array of deep learning-based techniques. These methods span a diverse spectrum, encompassing CNN [15], GANs [45, 25], and Transformer models [27, 28]. Their common goal is to directly translate the provided music into a continuous sequence of human poses. For instance, Esser et al. [8] encodes intricate visual elements into quantized patches, employing Transformers to generate contextually coherent images at high resolutions, bridging the visual and musical aspects of dance generation. In a different approach, Bailando [42] adopts VQ-VAE [48] to maintain temporal coherence across various music genres, a crucial element in dance generation, ensuring that the movements align with the music’s rhythm and style. And EDGE [47] leverages conditional diffusion models to craft human dance movements in direct response to musical cues, utilizing the potent audio feature extractor Jukebox [6].
@@ 3.4 Model
Our model architecture is depicted in Figure 2 and Figure 3. The inputs of our model are noise slice $z_{t}$, diffusion timestep $t$, music condition $c$, and beat condition $b$, where $z_{t}$, $c$, and $b$ have the same length. To ensure that the newly generated motion remains faithful to the preceding and future motions, we segment the entire noise sequence into smaller slices. Our model takes into consideration the previous motions and future noise distributions when generating the new motion slice. To elaborate on the process, we divide the complete noise sequence $z_{t}$ into $K$ slices at the timestep t, represented as $(z_{t1},z_{t2},...,z_{tK})$. Each noise slice $z_{tk}$ undergoes a cross-attention layer that considers the previously generated dance sequence $\hat{x}_{(k-1)}$ and the subsequent noise part $z_{t(k+1)}$ as shown:
$\hat{z}_{tk}=\text{Cross\_attention}(z_{tk},\text{cat}(\hat{x}_{(k-1)},z_{tk},z_{t(k+1)})).$ (7)
Each processed noise slice $\hat{z}_{tk}$ is subsequently sent into the decoder, which contains feature-wise linear modulation (FiLM) [39], alongside the segmented music feature $c_{k}$ and beat information $b_{k}$:
$\hat{x}_{k}=\text{Decoder}(\hat{z}_{tk},t,b_{k},c_{k}),$ (8)
the output is the generated dance slice $\hat{x}_{k}$.
@@
Figure 3: BADM processes noise slices in a bidirectional way. Generated dance slices are concatenated and sent to the local information decoder.
@@ 5 Discussion
In the final stage of our proposed model, we employ a local information decoder to enhance the harmony of the generated motions. We also experimented with a global information encoder to refine the output. In this process, the concatenated motions are processed through a transformer layer. However, the results were less than satisfying. The spatial freedom of the body’s hand and foot movements became constrained, and the body struggled to execute certain complex movements. Upon careful examination, we attribute this issue to the global information component, which compels the motion at each frame to account for actions in distant frames, ultimately restricting creative expression. These results further validate our decision to generate each motion based on nearby motions rather than considering the entire sequence of motions.

### v3

@@ Abstract
Dance serves as a powerful medium for expressing human emotions, but the lifelike generation of dance is still a considerable challenge. Recently, diffusion models have showcased remarkable generative abilities across various domains. They hold promise for human motion generation due to their adaptable many-to-many nature. Nonetheless, current diffusion-based motion generation models often create entire motion sequences directly and unidirectionally, lacking focus on the motion with local and bidirectional enhancement. When choreographing high-quality dance movements, people need to take into account not only the musical context but also the nearby music-aligned dance motions. To authentically capture human behavior, we propose a Bidirectional Autoregressive Diffusion Model (BADM) for music-to-dance generation, where a bidirectional encoder is built to enforce that the generated dance is harmonious in both the forward and backward directions. To make the generated dance motion smoother, a local information decoder is built for local motion enhancement. The proposed framework is able to generate new motions based on the input conditions and nearby motions, which foresees individual motion slices iteratively and consolidates all predictions. To further refine the synchronicity between the generated dance and the beat, the beat information is incorporated as an input to generate better music-aligned dance movements. Experimental results demonstrate that the proposed model achieves state-of-the-art performance compared to existing unidirectional approaches on the prominent benchmark for music-to-dance generation. Project page: https://czzhang179.github.io/badm.github.io/.
00footnotetext: This work was done during Canyu’s intership in PAII Inc.. 🖂Corresponding author.
@@ 1 Introduction
In response to the limitations inherent in current methodologies, we present the innovative bidirectional Autoregressive diffusion model (BADM). Our model introduces a novel strategy by breaking down the entire noise sequence into smaller, manageable slices. During the generation of each slice, our model considers the preceding dance sequences and forwarding noise distributions using a cross-attention layer in a bidirectional way. Then refined noise slices are sent into the decoder along with conditions iteratively. The output dance slices are then concatenated and refined by a local information decoder to ensure a cohesive and harmonious overall sequence from a local perspective.
@@
Figure 2: Our bidirectional Autoregressive diffusion model (BADM) employs a denoising mechanism to enhance dance sequences from time $t=T$ to $t=0$. BADM begins with a noisy sequence $z_{T}$ at time $T$, and proceeds to generate an estimated dance sequence $\hat{x}$. The denoising procedure is iteratively applied until $t=0$. Our autoregressive (AR) model treats the whole noise sequence as $K$ slices. On the left, we show our model processing the $k$-th slice at the diffusion timestep $t$. BADM is employed $K$ times within each BADM process.
@@ 2.2 Music-to-dance Generation
The challenge of generating dances that synchronize with the input music has fueled the development of a wide array of deep learning-based techniques. These methods span a diverse spectrum, encompassing CNN [16], GANs [47, 26], and Transformer models [28, 29]. Their common goal is to directly translate the provided music into a continuous sequence of human poses. For instance, Esser et al. [8] encodes intricate visual elements into quantized patches, employing Transformers to generate contextually coherent images at high resolutions, bridging the visual and musical aspects of dance generation. In a different approach, Bailando [44] adopts VQ-VAE [50] to maintain temporal coherence across various music genres, a crucial element in dance generation, ensuring that the movements align with the music’s rhythm and style. And EDGE [49] leverages conditional diffusion models to craft human dance movements in direct response to musical cues, utilizing the potent audio feature extractor Jukebox [6]. Diffusion based models also support human motion editing similar as inpainting [55, 54, 33, 10].
@@ 3.4 Model
Our model architecture is depicted in Figure 2 and Figure 3. The inputs of our model are noise slice $z_{t}$, diffusion timestep $t$, music condition $c$, and beat condition $b$, where $z_{t}$, $c$, and $b$ have the same length. To ensure that the newly generated motion remains faithful to the preceding and future motions, we segment the entire noise sequence into smaller slices. Our model takes into consideration the previous motions and future noise distributions when generating the new motion slice. To elaborate on the process, we divide the complete noise sequence $z_{t}$ into $K$ slices at the timestep t, represented as $(z_{t1},z_{t2},...,z_{tK})$. Each noise slice $z_{tk}$ undergoes a cross-attention layer that considers the previously generated dance sequence $\hat{x}_{(k-1)}$ and the subsequent noise part $z_{t(k+1)}$. Each processed noise slice $\hat{z}_{tk}$ is subsequently sent into the decoder, which contains feature-wise linear modulation (FiLM) [41], alongside the segmented music feature $c_{k}$ and beat information $b_{k}$, the output is the generated dance slice $\hat{x}_{k}$.
@@
Figure 3: BADM processes each noise slice $z_{k}$ in a bidirectional way. Generated dance slices are concatenated and sent to the local information decoder. We show this process at each timestep t.
@@ 5 Discussion
In the final stage of our proposed model, we employ a local information decoder to enhance the harmony of the generated motions. We also experimented with a global information encoder to refine the output. However, the results were less than satisfying. The spatial freedom of the body’s hand and foot movements became constrained, and the body struggled to execute certain complex movements. Upon careful examination, we attribute this issue to the global information component, which compels the motion at each frame to account for actions in distant frames, ultimately restricting creative expression. These results further validate our decision to generate each motion based on nearby motions rather than considering the entire sequence of motions.
@@ 7 Conclusion
Supplementary Material
FID results.
As claimed in EDGE, the FID metrics present limitations since AIST++ test set does not thoroughly cover the train distribution. Even so, we also present the FID results in Table 4. We use the codes from Bailando. The FACT and Bailando results were reported from their original papers. Our method gets the best results except for the $\text{FID}_{\text{g}}$ of Bailando, demonstrating its capability to generate smoother sequences. Bailando focuses on organizing the movements, and $\text{FID}_{\text{g}}$ metric reflects the quality of choreography.
Method $\text{FID}_{\text{k}}\downarrow$ $\text{FID}_{\text{g}}\downarrow$
GT 17.10 10.60
FACT 35.31 22.11
Bailando 28.16 9.62
EDGE 31.23 24.71
Ours 27.02 20.24
Table 4: FID results of the different methods.
The reason for using $z_{k+1}$ in AR encoder.
According to the characteristics of the DDPM diffusion model, as the timestep t increases, the added noise becomes increasingly subtle. Consequently, the noised sequence bears a greater resemblance to the original one. $z_{k+1}$ progressively offers clearer insights about the future sequence. Also, from the FID results in Table 5, it can be seen that the model using a bidirectional encoder yields smoother results compared to the model using a unidirectional encoder.
Method $\text{FID}_{\text{k}}\downarrow$ $\text{FID}_{\text{g}}\downarrow$
EDGE 31.23 24.71
w/o Beat 27.60 20.94
Unidirection 28.24 22.28
w/o LID 28.31 22.51
Ours 27.02 20.24
Table 5: FID results of the different settings.
Details about diffusion process.
We employ DDPM to elaborate on the entire diffusion process, with a diffusion timestep set to 1000. Throughout the training, we randomly select the time step t in each iteration, effectively constraining the training time. For inference, we leverage DDIM for expediting sampling.
Inference spreed.
In our setting, EDGE requires 9.54 seconds to generate a 5-second dance sequence, while our method takes 10.45 seconds. Although our method incurs a slightly higher time cost, the difference is marginal.